Which quantity in each of the following pairs is larger? 5.78×101km or 6.43×108cm?
1
Expert's answer
2019-08-26T05:48:27-0400
"1 km = 1000 m = 100000 cm (10^5 cm)"


"5.78*101 km = 5.78*101*10^5 cm = 583.78 * 10^5 cm = 5.8378*10^7 cm"


"6.43*108 cm = 694.44 cm = 6.9444*10^2cm"

Obviously, "5.78*101 km" is a larger numerical value.